# Litmus tests of the flat ${\Lambda\text{CDM}}$ model and model-independent measurement of $H_{0}r_{\text{d}}$ with LSST and DESI

### v1

@@ Appendix A Smooth Derivatives
$\hat{y}_{n+1}(t)$ $=\hat{y}_{n}(t)+\frac{\bm{\delta y}_{n}^{\intercal}\cdot\bm{C}^{-1}\cdot\bm{W}(t)}{\mathbbm{1}^{\intercal}\cdot\bm{C}^{-1}\cdot\bm{W}(t)},$ (A.1)
where $\mathbbm{1}^{\intercal}=(1,\cdots,1)$, the weight $\bm{W}$ and residual $\bm{\delta y}_{n}$ denote
$W_{i}(t)$ $=\exp\left[-\frac{(t-t_{i})^{2}}{2\Delta^{2}}\right]$ (A.2)
$\bm{\delta y}_{n}|_{i}$ $=y_{i}-\hat{y}_{n}(t_{i})$ (A.3)
and $\bm{C}^{-1}$ indicates the inverse of covariance matrix of the data. The smoothing width is set to $\Delta=0.3$ as mentioned in eq. (3.11).
@@ Acknowledgments
B. L. acknowledges the support of the National Research Foundation of Korea (NRF-2022R1F1A1076338 and RS-2023-00259422). AM thanks Richard Kessler and the LSST DESC TD team for the production of the LSST dataset.

### v3

@@ Appendix A Smooth Derivatives
Appendix B Error propagation
The errors are propagated using the usual formula: for $\bm{Y}=\bm{f}(\bm{X})$, the covariance $\bm{C}_{Y}$ is given by
$\bm{C}_{Y}$ $=\bm{J}^{\intercal}\cdot\bm{C}_{X}\cdot\bm{J},$ (B.1)
where $\bm{J}$ is the Jacobian of $\bm{f}$, and $\bm{Y}$ stands for c/$H_{0}r_{\text{d}}$, $\Theta$, or $\mathcal{O}_{k}$.
Since the correlation between two redshift bins in the BAO data is negligible, we can propagate the errors independently at fixed redshift $z$. We then have
$\bm{C}_{c/H_{0}r_{\text{d}}}$ $=\bm{J}_{c/H_{0}r_{\text{d}}}^{\intercal}\cdot\bm{C}_{\mathrm{BAO}}\cdot\bm{J}_{c/H_{0}r_{\text{d}}}$ (B.2)
where
$\bm{J}_{c/H_{0}r_{\text{d}}}$ $=\begin{pmatrix}\frac{1}{\mathcal{D}(z)}&0\\ 0&\frac{1}{\mathcal{D}^{\prime}(z)}\end{pmatrix}.$ (B.3)
For $\Theta$ and $\mathcal{O}_{k}$, the covariance matrix is simply the variance. The variance of $\Theta$ is obtained by
$\sigma_{\Theta}^{2}$ $=\bm{J}_{\Theta}^{\intercal}\cdot\bm{C}_{\text{BAO}}\cdot\bm{J}_{\Theta},$ (B.4)
where
$\bm{J}_{\Theta}$ $=\Theta\begin{pmatrix}\frac{r_{\text{d}}}{d_{\text{M}}}\\ \frac{r_{\text{d}}}{d_{\text{H}}}\end{pmatrix},$ (B.5)
and that of $\mathcal{O}_{k}$ by
$\sigma_{\mathcal{O}_{k}}=2\frac{\Theta}{\mathcal{D}^{2}}\sigma_{\Theta}.$ (B.6)
In all cases, the total $\chi^{2}$ for reconstruction $n$ is obtained by
$\chi^{2}_{n}=\chi^{2}_{\text{SN},n}+\chi^{2}_{\text{BAO},n}.$ (B.7)
@@ Acknowledgments
The authors thank Kushal Loda for constructive discussions. B. L. acknowledges the support of the National Research Foundation of Korea (NRF-2022R1F1A1076338 and RS-2023-00259422). A.M. thanks Richard Kessler and the LSST DESC TD team for the production of the LSST dataset. A. S. would like to acknowledge the support by National Research Foundation of Korea NRF2021M3F7A1082056 and the support of the Korea Institute for Advanced Study (KIAS) grant funded by the government of Korea.